# Robust $Q$-Learning for finite ambiguity sets

### v2

@@ 6.2. Proofs of the results from Section 2, Section 3, and Section 4
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)~{}\middle|~{}\mathcal{G}_{t}\right]\right|$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right]\right|$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\Bigg{|}\min_{i=1,...,N}\left(\mathbb{E}_{\mathbb{P}^{(i)}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]\right)-Q^{*}(x,a)\Bigg{|}$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\Bigg{|}\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q_{t}(X_{1},b)\right]-Q^{*}(x,a)\Bigg{|}.$
By using the definition of the operator $\mathcal{H}$ from (6.2) we therefore obtain
$|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}|\mathcal{H}Q_{t}(x,a)-Q^{*}(x,a)|.$
@@
(6.6) $|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}|\mathcal{H}Q_{t}(x,a)-\mathcal{H}Q^{*}(x,a)|.$
@@
$\|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(\cdot,\cdot)|\mathcal{G}_{t}]\|_{\infty}\leq\|\mathcal{H}Q_{t}-\mathcal{H}Q^{*}\|_{\infty}\leq\alpha\|Q_{t}-Q^{*}\|_{\infty}=\alpha\|\Delta_{t}\|_{\infty}$
@@
(6.7) $\operatorname{Var}$ (${}_{\mathbb{P}_{x_{0},\textbf{a}}}(F_{t}(x,a)|\mathcal{G}_{t})$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\cdot\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right)$
$\leq\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right)\right)$
@@
$=\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})\right).$
We define $C_{r}:=\max_{y_{0},y_{1}\in\mathcal{X},\ b\in A}|r(y_{0},b,y_{1})|<\infty$ and compute as an upper bound for the integrand from (6.7) that
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})$
$\leq C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\max_{b^{\prime}\in A}Q^{*}(X_{t+1},b^{\prime})$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}(\max_{b^{\prime}\in A}Q^{*}(y,b^{\prime})-\max_{b\in A}Q_{t}(y,b))$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}\max_{b\in A}|Q^{*}(y,b)-Q_{t}(y,b)|$
@@
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}+\alpha(\min_{y^{\prime}\in\mathcal{X}}\min_{b\in A}(Q^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)))$
$\geq-C_{r}+\alpha(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)|)$
$\geq-C_{r}+\alpha\min_{y\in\mathcal{X}}(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q^{*}(y^{\prime},b)-Q_{t}(y,b)|)$
$\geq-C_{r}-\alpha(\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q_{t}(y,b)-Q^{*}(y,b)|+|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|)$
$\geq-C_{r}-\alpha\|\Delta_{t}\|_{\infty}-\alpha\max_{(y^{\prime},y)\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|=:m\in\mathbb{R}.$
@@
$\leq$ $\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})\right)$
@@
$\leq$ $\frac{1}{4}\Big{(}C_{r}+\alpha\|\Delta_{t}\|_{\infty}+C_{r}+\alpha\|\Delta_{t}\|_{\infty}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}$
$=$ $\frac{1}{4}\Big{(}C_{r}+C_{r}+2\alpha\|\Delta_{t}\|_{\infty}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}$
$\leq$ $\frac{1}{2}\left(\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\left(\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\Bigg{(}\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}$
$+4\alpha^{2}+\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}\|\Delta_{t}\|_{\infty}^{2}\Bigg{)}$
$\leq$ $\left(4\alpha^{2}+\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}\right)(1+\|\Delta_{t}\|_{\infty}^{2})$
$\leq$ $C\cdot(1+\|\Delta_{t}\|_{\infty})^{2},$
@@
(6.8) $C:=\left(4\alpha^{2}+(2C_{r}+\alpha\max_{(y^{\prime},y)\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|)^{2}\right).$

### v3

@@ 7.2. Proofs of the results from Section 2, Section 3, and Section 4
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)~\middle|~\mathcal{G}_{t}\right]\right|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right]\right|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\widetilde{\mathcal{H}}_{t}Q_{t}(x,a)-Q_{t}^{*}(x,a)\right|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\mathcal{H}_{t}Q_{t}(x,a)-Q_{t}^{*}(x,a)+\widetilde{\mathcal{H}}_{t}Q_{t}(x,a)-{\mathcal{H}}_{t}Q_{t}(x,a)\right|$
$\leq 1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left(\left|\mathcal{H}_{t}Q_{t}(x,a)-Q_{t}^{*}(x,a)\right|+c_{t}\right)$
@@
(7.15) $|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|\leq 1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left(|\mathcal{H}_{t}Q_{t}(x,a)-\mathcal{H}_{t}Q_{t}^{*}(x,a)|+c_{t}\right).$
@@
$\|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(\cdot,\cdot)|\mathcal{G}_{t}]\|_{\infty}\leq\|\mathcal{H}_{t}Q_{t}-\mathcal{H}_{t}Q_{t}^{*}\|_{\infty}+c_{t}\leq\alpha\|Q_{t}-Q_{t}^{*}\|_{\infty}+c_{t}=\alpha\|\Delta_{t}\|_{\infty}+c_{t}$
@@
[…]
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\cdot\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right)$
$\leq\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right)\right)$
@@
$=\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})\right).$
We recall $C_{r}:=\max_{y_{0},y_{1}\in\mathcal{X},\ b\in A}|r(y_{0},b,y_{1})|<\infty$ and compute as an upper bound for the integrand from (7.16) that
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})$
$\leq C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\max_{b^{\prime}\in A}Q_{t}^{*}(X_{t+1},b^{\prime})$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}(\max_{b^{\prime}\in A}Q_{t}^{*}(y,b^{\prime})-\max_{b\in A}Q_{t}(y,b))$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}\max_{b\in A}|Q_{t}^{*}(y,b)-Q_{t}(y,b)|$
@@
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}+\alpha(\min_{y^{\prime}\in\mathcal{X}}\min_{b\in A}(Q_{t}^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)))$
$\geq-C_{r}+\alpha(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q_{t}^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)|)$
$\geq-C_{r}+\alpha\min_{y\in\mathcal{X}}(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q_{t}^{*}(y^{\prime},b)-Q_{t}(y,b)|)$
$\geq-C_{r}-\alpha(\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q_{t}(y,b)-Q_{t}^{*}(y,b)|+|Q_{t}^{*}(y,b)-Q_{t}^{*}(y^{\prime},b)|)$
$\geq-C_{r}-\alpha\|\Delta_{t}\|_{\infty}-\alpha\max_{(y^{\prime},y)\in\mathcal{X},\ b\in A}|Q_{t}^{*}(y,b)-Q_{t}^{*}(y^{\prime},b)|$
$\geq-C_{r}-\alpha\|\Delta_{t}\|_{\infty}-\alpha\frac{2C_{r}}{1-\alpha}=:m\in\mathbb{R}.$
@@
$\leq$ $\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})\right)$
@@
$\leq$ $\frac{1}{4}\Big(C_{r}+\alpha\|\Delta_{t}\|_{\infty}+C_{r}+\alpha\|\Delta_{t}\|_{\infty}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}$
$=$ $\frac{1}{4}\Big(C_{r}+C_{r}+2\alpha\|\Delta_{t}\|_{\infty}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}$
$\leq$ $\frac{1}{2}\left(\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\left(\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\Bigg(\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}+4\alpha^{2}+\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}\|\Delta_{t}\|_{\infty}^{2}\Bigg)$
$\leq$ $\left(4\alpha^{2}+\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}\right)(1+\|\Delta_{t}\|_{\infty}^{2})\leq C\cdot(1+\|\Delta_{t}\|_{\infty})^{2},$
@@
(7.17) $C:=\left(4\alpha^{2}+(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha})^{2}\right).$